OpenStudy (anonymous):

Find the exact value of the trigonometric function. sin 315°

OpenStudy (anonymous):

isnt the answer \[-1/\sqrt{2}\]

OpenStudy (anonymous):

or \[-1\sqrt{2}/2\]

OpenStudy (anonymous):

-rt(2)/2

OpenStudy (anonymous):

-rt(2)/2??? im lost. im sorry my teacher never said anything about that

OpenStudy (anonymous):

same as ur ans ....

OpenStudy (anonymous):

−1/√2

OpenStudy (anonymous):

??? that

OpenStudy (anonymous):



OpenStudy (anonymous):

ok yeah thats what i thought. the sign can only be on the top right? its never suppose to be on the bottom.

OpenStudy (anonymous):

yaa... hope u r clear ...

OpenStudy (anonymous):

yeah that makes so much sense.

OpenStudy (anonymous):

thank you darling.

OpenStudy (anonymous):

4th quadrant so sin is neg. this is 45 deg angle so \[\sin 315^{\circ} = -\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

top, bottom, middle is all the same... just a negative number

OpenStudy (anonymous):

315 = 360-45 then sine315=-sine 45=-1/√2